OpenStudy (anonymous):

Find the slope of the line that passes through (3, -4) and (-1, 2).

OpenStudy (anonymous):

-3/2

OpenStudy (hoblos):

slope = variation of y over variation of x slope= (2+4)/(-1-3) = 6/(-4) = -3/2